OpenStudy (anonymous):

Let \( f \) be a function in two variables defined for natural numbers and satisfying 1) \( f(x, x) = x \) 2)\( f(x, y) = f(y, x) \) 3)\( (x + y)f(x, y) = y.f (x, x + y) \) Then find the value of \( f(252, 90) \).

OpenStudy (turingtest):

f=0 works but I suppose you wanted a non-trivial answer...

OpenStudy (binary3i):

is it 19/5

OpenStudy (binary3i):

sorry , half the way

OpenStudy (binary3i):

i dont know but is it 5

OpenStudy (anonymous):

The answer is not 5 or 19/5.

OpenStudy (anonymous):

2520?

OpenStudy (anonymous):

1260 is the answer. LCM ;)

OpenStudy (anonymous):

yay! joemath

OpenStudy (anonymous):

Slight variation, what would the answer be if the third rule was: \[3) f(x,y)=f(x,y-x)\]

OpenStudy (anonymous):

sorry Professor Joemath

OpenStudy (anonymous):

wait, miss type >.< it should be y+x in the second argument....now that i think about it lol

OpenStudy (anonymous):

So third rule:\[f(x,y) = f(x,y+x)\]

OpenStudy (anonymous):

GCD isn't ?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

yep :) these were interesting problems. Writing out proofs for the solutions were fun.

OpenStudy (anonymous):

but foolformath's third rule makes it hard

OpenStudy (anonymous):

for me

OpenStudy (anonymous):

If you try it with smaller numbers its easier to see.

OpenStudy (anonymous):

no really Ishaan, you just don't know (yet) how to tackle these kinds of problems :)

OpenStudy (anonymous):

joe is professor ??

OpenStudy (anonymous):

im not >.> akshay started calling me that lol. Im a lowly undergraduate.

OpenStudy (anonymous):

lol cool man :D

OpenStudy (anonymous):

btw joe what about \[ f(x,y)=f(x,y-x) \] ?

OpenStudy (anonymous):

ah that was a mistype. although it would lead to the same solution, GCD.

OpenStudy (anonymous):

yeah :D